Data on the percentage of 200 hotels in each of the three large cities across the world on whether minibar charges are correctly posted at checkout are given below. At the 0.05 level of significance, you want to know if there is evidence of a difference in the proportion of hotels that correctly post minibar charges among the three cities.

HONG KON NEW YORK PARIS
YES 86% 76% 78%
NO 14% 24% 22%

9) Referring to Scenario 6, what are the values of the absolute difference and the critical range for the Marascuilo procedure to test for the difference in proportions between Hong Kong and Paris using a 0.05 level of significance?
Answer #1

9) The absolute difference is : p1 - p3 = (0.86 - 0.78) = 0.08 . where p1 = proportion of 'Yes' for Hongkong and p3 = proportion of 'Yes' for Paris.

The formula for the critical range is :

here, k=number of different proportions in the data = 3; and n1 = n3 = 200;

So,